# Treewidth 2 in the Planar Graph Product Structure Theorem

### v2

@@ 1 Introduction
Graph product structure theory describes graphs in complicated graph classes as subgraphs of products of graphs in simpler graph classes, typically with bounded treewidth or bounded pathwidth. As defined in Section 2, the treewidth of a graph $G$, denoted by $\operatorname{tw}(G)$, is the standard measure of how similar $G$ is to a tree.
As illustrated in Figure 1, the strong product $A\mathbin{\boxtimes}B$ of graphs $A$ and $B$ has vertex-set
@@
$V(A)\times V(B)$, where distinct vertices $(v,x),(w,y)$ are adjacent if:
@@ 3 Proof of Theorem 2
We may assume that $G$ has an edge $ab$ such that $G^{\prime}:=G-a-b$ is planar. By Corollaries 7 and 4, $G^{\prime}$ has a matching $M^{\prime}$ such that $(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$ for some apex-forests $H_{\_}1$ and $H_{\_}2$. Thus, for the matching $M:=M^{\prime}\cup\{ab\}$ we have $G/M\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$. Hence $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$. Here we use the fact that if $M$ is a matching in a graph $G$, then $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G/M)\boxtimes K_{\_}2$. ∎
@@ 4 Proof of Theorem 3
The next result follows from Lemmas 12 and 9, which implies Theorem 3.

### v3

@@ 1 Introduction
Graph product structure theory describes graphs in complicated graph classes as subgraphs of products of graphs in simpler graph classes, typically with bounded treewidth or bounded pathwidth. As defined in Section 2, the treewidth of a graph $G$, denoted by $\operatorname{tw}(G)$, is the standard measure of how similar $G$ is to a tree. As illustrated in Figure 1, the strong product $A\mathbin{\boxtimes}B$ of graphs $A$ and $B$ has vertex-set $V(A)\times V(B)$,
@@
where distinct vertices $(v,x),(w,y)$ are adjacent if:
@@ 3 Proof of Theorem 2
We may assume that $G$ has an edge $ab$ such that $G^{\prime}:=G-a-b$ is planar. By Corollaries 1 and 4, $G^{\prime}$ has a matching $M^{\prime}$ such that $(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$ for some apex-forests $H_{\_}1$ and $H_{\_}2$. Thus, for the matching $M:=M^{\prime}\cup\{ab\}$ we have $G/M\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$. Hence $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$. Here we use the fact that if $M$ is a matching in a graph $G$, then $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G/M)\boxtimes K_{\_}2$. ∎
@@ 4 Proof of Theorem 3
The next result follows from Lemmas 10 and 1, which implies Theorem 3.
@@
Acknowledgements
Distel is supported by an Australian Government Research Training Program Scholarship. Hendrey and Wood are supported by the Australian Research Council. Karol and Yip are supported by Monash Graduate Scholarships.